ctn 30 degrees, writing the exact value.
Answer by kietra(57)

To find the ctn 30 degrees, draw a triangle on the unit circle with the inside angle 30 degrees. On the unit circle, it has a right angle and we know that cotangent = adjacent side/opposide side. (Do you know SOH-CAH-TOA where tangent is opposite over adjacent and cot is inverse of tangent so adjacent over opposite?) Now, in the unit triangle, the side opposite the 30 degree angle is always 1, the side opposite the 60 degree angle is always 1* sqrt(3) and the hypotenuse is always 2. So, cot30=a/o=sqrt(3)/1.
cot30=sqrt(3)